A jewler has two bars of gold alloy in stock, one 12 carets and the other 18 carters (24 carat gold is pure gold, 12 carat gold is 12/24 pure, 18-carat gold is 18/24 pure and so on). How many grams of each alloy must be mixed to obtain 10 grams of 14-carat gold?
Answer by mananth(16946)

percent ----------------quantity
Gold 12 carat 12----------------x gms
Gold 18 carat 18----------------10-x gms
Total 14.00% ----------------10 gms

12x+18(10-x)=14*10
12x+180-18x=140
12x-18x =140-180
-6x= -40
/-6
x=6.67gms 12 carat
3.33 gms18 carat